# Control of seizure-like dynamics in neuronal populations with excitability adaptation related to ketogenic diet

### v1

@@ I Introduction
The recent advent of the next-generation neural mass (NGNM) models [1, 2, 3, 4, 5] has allowed for a deeper understanding of multistability and critical transitions in the dynamics of neuronal populations, and has provided us with the means to systematically explore the control strategies for switching between the states interpreted in terms of neurophysiology as regular or pathological ones. Earlier phenomenological neural mass models [6, 7, 8] described the dynamics of neuronal populations in terms of firing rate equations, typically assuming a sigmoid-like transfer function between the mean membrane potential and the mean firing rate. Such models relied on the assumption of uncorrelated local dynamics and could neither capture the scenarios where the state of synchrony undergoes qualitative change nor establish an exact relationship between the features of local and collective dynamics. These issues have been resolved in the NGNM models, based on applying the Ott-Antonsen [9, 10, 11] or the equivalent Lorentzian Ansatz [1] to populations of theta neurons and quadratic integrate-and-fire neurons, respectively. NGNM models characterize population dynamics in terms of mean-field equations for the mean firing rate and the mean membrane potential. While the notion of spike synchrony, classically invoked in neuroscience, is reflected in the increase of the mean firing rate, the NGNM framework further establishes a conformal mapping [1] between the mean firing rate and the population synchrony described by the Kuramoto order parameter, a synchrony measure inherited from the theory of coupled phase oscillators. Given that the mapping is a real valued function of the complex argument, different values of the Kuramoto order parameter may correspond to the same mean firing rate. NGNM models have the advantage of being exact in the thermodynamic limit and are amenable to bifurcation analysis. So far, NGNM models have been used to address a broad range of theoretical problems, including understanding the impact and interplay of chemical and electrical synapses [12, 13, 14], the effects of deterministic or stochastic drive [15, 16, 17, 18, 19], presence of quenched randomness in network topology [20], finite-size fluctuations [21, 22], interaction of neuronal populations [23, 24, 25], as well as formation of bumps and waves in neural fields [26, 27, 28]. Moreover, a wide spectrum of applications has been considered, from explaining the onset of gamma and theta-nested gamma oscillations [29, 30, 31, 32, 33], cross-frequency couplings [34], abnormal beta rebound in schizophrenia [35], the impact of deep brain stimulation [36], to studies on working memory [24, 37], propagation of epileptic seizures [38], and modeling of the entire brain functional networks [4].
@@
The paper is organized as follows. In Sec. II, we introduce the model of a globally coupled heterogeneous population of intrinsically excitable or tonic spiking excitatory quadratic integrate-and-fire neurons supplied by the local terms for the ATP-dependent hyperpolarizing currents and an equation for the global dynamics of ATP concentration. Using the approach from [1], in Sec. III we derive the corresponding mean-field model, which compared to the classical results, contains additional dependencies of the mean firing rate and mean membrane potential on the ATP concentration level. Section IV provides the bifurcation analysis of the reduced system, revealing two qualitatively different bifurcation scenarios in terms of transitions to synchrony and the system’s multistability depending on the coupling strength. In Sec. V we demonstrate mechanisms for controlled switching between the homeostatic and seizure-like states, including external stimulation, parametric and dynamical perturbation. Section VI provides a summary and discussion of our results.
@@ II Model of microscopic dynamics
$S(t)=\frac{1}{N}\sum_{j=1}^{N}\sum_{k\backslash t_{j}^{k}<t}\delta(t-t_{j}^{k}),$ (3)
where $t_{j}^{k}$ is the time of the k-th spike of neuron $j$. The excitability adaptation triggered by the changes in the energy metabolism due to the ketogenic diet is mediated by the ATP-dependent hyperpolarizing currents $I_{j}^{\mathrm{ATP}}$. The latter reflects the action of the ATP-gated potassium channels in neuronal membranes activated by the diet [69, 74, 75], and are assumed to have a uniform conductance $\alpha$. Note that the gating of the potassium channels is in fact a complex process governed by the exchange of sodium and ATP [71, 70]. Nevertheless, for simplicity, here we adopt the model for $I_{j}^{\mathrm{ATP}}$ from [69] which omits the detailed equation for the sodium dynamics. To close the system  (1), one requires an additional equation for the dynamics of ATP concentration $C(t)$
@@
which should be complemented with Eq. (4) for the dynamics of ATP concentration. The spiking events from the representation  (1) just correspond to the phase variables’ crossing the $\pi$ value. The Lorentzian distribution of local bifurcation parameters $\eta_{j}$ is generated deterministically via $\eta_{j}=\bar{\eta}+\Delta\tan((\pi/2)\frac{2j-N-1}{N+1})$ [12]. The numerical simulations of the model  (5), (4) are carried out for the fixed parameter set $\Delta=1.0,\alpha=1.0,\epsilon=1.0,\tilde{C}=1.0,N=10^{4}$ and varying $\bar{\eta}$, $K$ and $\tau$. While $\bar{\eta}$ and $K$ respectively characterize the intrinsic population dynamics and the impact of interactions in modifying the local dynamics, $\tau$ may be seen as the main parameter that describes the metabolic changes associated with the ketogenic diet. For numerical integration of the model  (5) and  (4), we use an error-controlled 5/4 Runge-Kutta scheme called Tsit5 implemented in the Julia package DifferentialEquations.jl [78]. During the simulation, the impact of the spikes and the ATP consumption are handled separately. In particular, we determine $S(t)$ by counting the number of spikes within a given time step and then apply changes to the phases and the ATP concentration according to $\theta_{j}(t)^{+}=\theta_{j}(t)^{-}+(1+\cos(\theta_{j}(t)^{-}))KS(t)$ and $C(t)^{+}=C(t)^{-}-S(t)\epsilon C(t)^{-}/\tilde{C}$, where the superscripts -- and $+$ denote the values before and after the impact of the spikes, respectively. In a subsequent comparison with the mean-field model, we present the mean firing rate of the population of theta neurons as a time-smoothed signal of $S(t)$ using a rectangular convolution function with the time width $0.1$, similar to [12].
@@ III Mean-field system
Applying the general reduction approach from [1] to the microscopic model  (1)- (4), in this Section we derive the mean-field system describing the collective behavior of a neuronal population affected by the excitability adaptation reflecting the metabolic changes related to ketogenic diet. In the thermodynamic limit $N\rightarrow\infty$, the population state can be characterized by the conditional density function $\rho(V|\eta,t)$, such that $\rho(V|\eta,t)dV$ gives the fraction of neurons with membrane potential within the interval $(V,V+dV)$ and local bifurcation parameter $\eta$ at time $t$. The density function satisfies the continuity equation
@@ V Control of switching dynamics
Figure 7: Switches between the coexisting seizure-like and homeostatic states triggered by sudden changes in ATP level, indicated by arrows in (c). A positive ATP shock of finite duration induces the switch from the homeostatic to the seizure-like state. The homeostatic state is regained following an instantaneous, small-amplitude negative ATP shock to the seizure-like state. (b)-(d) Comparison between the time traces $r(t)$, $v(t)$ and $C(t)$ for the mean-field model (11) (dashed lines) and the corresponding variables of the microscopic system (5) (solid lines); (e) spike raster plot for (5). Parameters: $\tau=2.9$, $K=15.0$, $\bar{\eta}={-}2.7$, $\Delta=\epsilon=\alpha=\tilde{C}=1.0$ (diamond in Fig. 2(b)) and $N=10^{4}$.
@@ VI Summary and discussion
Within the last two decades, there has been a growing awareness that the feedback of neuronal activity with energy metabolism plays a pivotal role both in maintaining the neuronal homeostasis [56, 86, 57, 58, 87, 49, 42] and in development of certain neurological disorders, such as epilepsy [52, 53, 54, 55] or Parkinson’s disease [64, 65]. Nevertheless, many earlier computational studies have suffered from improperly including the energy constraints, using oversimplified models of local dynamics or implementing too small networks with rather strong finite-size effects. The recent theoretical breakthrough associated with the advent of NGNM models has provided us with a fascinating ability to address the mechanisms of neurological disorders with a strong metabolic component, and to even gain a foothold in understanding the impact of certain therapeutic procedures. In the present paper, we have introduced and analyzed a model of a heterogeneous, globally coupled population of quadratic integrate-and-fire neurons where the local excitability is modified via an ATP-sensitive potassium current. Implementing the classical reduction approach from [1], we have derived a three-dimensional mean-field model of population dynamics. The model is intended as a first step toward theoretically explaining the mechanism and apparent effectiveness of ketogenic diet, a long-standing treatment for epilepsy, which nowadays receives a revived interest for helping patients with a drug-resistant form of the disease [50, 52].
@@
We believe that our results, together with [38, 35], have set the stage for opening a new direction of research aimed at theoretical understanding of neurological disorders and associated therapeutic procedures by implementing the next-generation neural mass models. This is expected to provide a valuable complement to experimental studies in at least two ways. Firstly, it will allow for a qualitative insight into multistability which may be lacking or not be easily accessible to experiments; and secondly, it may help in developing effective strategies to control the switching dynamics, especially in terms of suppressing the pathological states. Concerning the research on the ketogenic diet, the next step could involve incorporating some other ingredients that may be related to its therapeutic effects. In particular, one may attempt to derive a “compartmental” model of local dynamics that could distinguish between the two different ATP production processes, namely the classical, glycolytic one near the membrane and the diet-related one unfolding in mitochondria. Also, one may attempt to circumvent the limitations of the current model associated with the assumed uniformity of ATP concentration over the population, aiming to accommodate for the neuron-specific ATP concentration variables that depend on the spike trains of individual neurons [69]. The microscopic system introduced in the latter way would likely show significant deviations from the current mean-field model, similar to those observed in [37, 40]. Finally, the immediate impact of ketone bodies on inhibition of excitatory synaptic channels could be taken into account as a factor modulating the synaptic efficacy. More generally, an interesting challenge would be to understand how the classical physical picture developed for activity of neuronal populations within the NGNM framework changes by incorporating the different forms of homeostatic plasticity [94], which involve negative feedback loops to constrain the network activity within the desired physiological limit.

### v2

@@ I Introduction
The recent advent of the next-generation neural mass (NGNM) models [1, 2, 3, 4, 5] has allowed for a deeper understanding of multistability and critical transitions in the dynamics of neuronal populations, and has provided us with the means to systematically explore the control strategies for switching between the states interpreted in terms of neurophysiology as regular or pathological ones. Earlier phenomenological neural mass models [6, 7, 8] described the dynamics of neuronal populations in terms of firing rate equations, typically assuming a sigmoid-like transfer function between the mean membrane potential and the mean firing rate. Such models relied on the assumption of uncorrelated local dynamics and could neither capture the scenarios where the state of synchrony undergoes qualitative change nor establish an exact relationship between the features of local and collective dynamics. These issues have been resolved in the NGNM models, based on applying the Ott-Antonsen [9, 10, 11] or the equivalent Lorentzian Ansatz [1] to populations of theta neurons and quadratic integrate-and-fire neurons, respectively. NGNM models characterize population dynamics in terms of mean-field equations for the mean firing rate and the mean membrane potential. While the notion of spike synchrony, classically invoked in neuroscience, is reflected in oscillations of the mean firing rate [12], the NGNM framework further establishes a relation between the mean firing rate and the population synchrony described by the complex Kuramoto order parameter, a synchrony measure inherited from the theory of coupled phase oscillators [1, 4]. NGNM models have the advantage of being exact in the thermodynamic limit and are amenable to bifurcation analysis. So far, NGNM models have been used to address a broad range of theoretical problems, including understanding the impact and interplay of chemical and electrical synapses [13, 14, 15], the effects of deterministic or stochastic drive [16, 17, 18, 19, 20], presence of quenched randomness in network topology [21], finite-size fluctuations [22, 23], interaction of neuronal populations [24, 25, 26], as well as formation of bumps and waves in neural fields [27, 28, 29]. Moreover, a wide spectrum of applications has been considered, from explaining the onset of gamma and theta-nested gamma oscillations [12, 30, 31, 32, 33], cross-frequency couplings [34], abnormal beta rebound in schizophrenia [35], the impact of deep brain stimulation [36], to studies on working memory [25, 37], propagation of epileptic seizures [38], and modeling of the entire brain functional networks [4].
@@
The paper is organized as follows. In Sec. II, we introduce the model of a globally coupled heterogeneous population of intrinsically excitable or tonic spiking excitatory quadratic integrate-and-fire neurons supplied by the individual terms for the ATP-dependent hyperpolarizing currents and an equation for the global dynamics of ATP concentration. Using the approach from [1], in Sec. III we derive the corresponding mean-field model, which compared to the classical results, contains additional dependencies of the mean firing rate and mean membrane potential on the ATP concentration level. Section IV provides the bifurcation analysis of the reduced system, revealing two qualitatively different bifurcation scenarios in terms of transitions to synchrony and the system’s multistability depending on the coupling strength. In Sec. V we demonstrate mechanisms for controlled switching between the homeostatic and seizure-like states, including external stimulation, parametric and dynamical perturbation. Section VI provides a summary and discussion of our results.
@@ II Model of microscopic dynamics
$S(t)=\frac{1}{N}\sum_{j=1}^{N}\sum_{k\backslash t_{j}^{k}\leq t}\delta(t-t_{j}^{k}),$ (3)
where $t_{j}^{k}$ is the time of the k-th spike of neuron $j$.
The excitability adaptation triggered by the changes in the energy metabolism due to the ketogenic diet is mediated by the ATP-dependent hyperpolarizing currents $I_{j}^{\mathrm{ATP}}$, which may be regarded as an effective parametric perturbation to the local bifurcation parameters $\eta_{j}$. The hyperpolarizing currents reflect the action of the ATP-gated potassium channels in neuronal membranes activated by the diet [69, 74, 75, 50, 76], and are assumed to have a uniform conductance $\alpha$. Note that the gating of the potassium channels is in fact a complex process governed by the exchange of sodium and ATP [71, 70]. Nevertheless, for simplicity, here we adopt the model for $I_{j}^{\mathrm{ATP}}$ from [69] which omits the detailed equation for the sodium dynamics. To close the system  (1), one requires an additional equation for the dynamics of ATP concentration $C(t)$
@@
which should be complemented with Eq. (4) for the dynamics of ATP concentration. The spiking events from the representation  (1) just correspond to the phase variables’ crossing the $\pi$ value. The Lorentzian distribution of local bifurcation parameters $\eta_{j}$ is generated deterministically via $\eta_{j}=\bar{\eta}+\Delta\tan((\pi/2)\frac{2j-N-1}{N+1})$ [13]. The numerical simulations of the model  (5), (4) are carried out for the fixed parameter set $\Delta=1.0,\alpha=1.0,\epsilon=1.0,\tilde{C}=1.0,N=10^{4}$ and varying $\bar{\eta}$, $K$ and $\tau$. While $\bar{\eta}$ and $K$ respectively characterize the intrinsic population dynamics and the impact of interactions in modifying the local dynamics, $\tau$ may be seen as the main parameter that describes the metabolic changes associated with the ketogenic diet, consistent with both in-vivo and in-vitro findings that the diet affects the ATP production capacity [50, 79, 80, 81, 82, 83]. For numerical integration of the model  (5) and  (4), we use an error-controlled 5/4 Runge-Kutta scheme called Tsit5 implemented in the Julia package DifferentialEquations.jl [84]. During the simulation, the impact of the spikes and the ATP consumption are handled separately. In particular, we determine $S(t)$ by counting the number of spikes within a given time step and then apply changes to the phases and the ATP concentration according to $\theta_{j}(t)^{+}=\theta_{j}(t)^{-}+(1+\cos(\theta_{j}(t)^{-}))KS(t)$ and $C(t)^{+}=C(t)^{-}-S(t)\epsilon C(t)^{-}/\tilde{C}$, where the superscripts -- and $+$ denote the values before and after the impact of the spikes, respectively. In a subsequent comparison with the mean-field model, we present the mean firing rate of the population of theta neurons as a time-smoothed signal of $S(t)$ using a rectangular convolution function with the time width $0.1$, similar to [13].
@@ III Mean-field system
Applying the general reduction approach from [1] to the microscopic model  (1)– (4), in this Section we derive the mean-field system describing the collective behavior of a neuronal population affected by the excitability adaptation reflecting the metabolic changes related to ketogenic diet. In the thermodynamic limit $N\rightarrow\infty$, the population state can be characterized by the conditional density function $\rho(V|\eta,t)$, such that $\rho(V|\eta,t)dV$ gives the fraction of neurons with membrane potential within the interval $(V,V+dV)$ and local bifurcation parameter $\eta$ at time $t$. The density function satisfies the continuity equation
@@ V Control of switching dynamics
Figure 7: Switches between the coexisting seizure-like and homeostatic states triggered by sudden changes in ATP level, indicated by arrows in (c). A positive ATP shock of finite duration induces the switch from the homeostatic to the seizure-like state. The homeostatic state is regained following an instantaneous, small-amplitude negative ATP shock to the seizure-like state. (a)-(c) Comparison between the time traces $r(t)$, $v(t)$ and $C(t)$ for the mean-field model (11) (dashed lines) and the corresponding variables of the microscopic system (5) (solid lines); (d) spike raster plot for (5). Parameters: $\tau=2.9$, $K=15.0$, $\bar{\eta}={-}2.7$, $\Delta=\epsilon=\alpha=\tilde{C}=1.0$ (diamond in Fig. 2(b)) and $N=10^{4}$.
@@
The examples provided in Sec. V indicate a spectrum of strategies that may in general be applied to regain the homeostatic state, and selecting an optimal strategy in any particular instance is dependent on the corresponding system parameters.
@@ VI Summary and discussion
Within the last two decades, there has been a growing awareness that the feedback of neuronal activity with energy metabolism plays a pivotal role both in maintaining the neuronal homeostasis [56, 92, 57, 58, 93, 49, 42] and in development of certain neurological disorders, such as epilepsy [52, 53, 54, 55] or Parkinson’s disease [64, 65]. Nevertheless, many earlier computational studies have suffered from improperly including the energy constraints, using oversimplified models of local dynamics or implementing too small networks with rather strong finite-size effects. The recent theoretical breakthrough associated with the advent of NGNM models has provided us with a fascinating ability to address the mechanisms of neurological disorders with a strong metabolic component, and to even gain a foothold in understanding the impact of certain therapeutic procedures. In the present paper, we have introduced and analyzed a model of a heterogeneous, globally coupled population of quadratic integrate-and-fire neurons where the individual excitability is modified via an ATP-sensitive potassium current. Implementing the classical reduction approach from [1], we have derived a three-dimensional mean-field model of population dynamics. The model is intended as a first step toward theoretically explaining the mechanism and apparent effectiveness of ketogenic diet, a long-standing treatment for epilepsy, which nowadays receives a revived interest for helping patients with a drug-resistant form of the disease [50, 52].
@@
One should point out that the current model does not provide an exhaustive description of the effects of epilepsy and ketogenic diet on the local neuronal kinetics. In particular, epilepsy is known to impact the dynamics of leaky sodium, potassium and chloride currents by triggering changes in the corresponding reversal potentials [100, 101, 102]. Also, the current model is limited by the fact that it does not consider the reversal potential of the ATP-dependent potassium current and the way in which it is affected by the ketogenic diet. While this is partly related to the lack of conclusive experimental evidence, it is also associated with the fact that these changes likely cannot be considered isolated from the dynamics of the leaky currents, including their implicit dependencies on surrounding glia, whose taking into account would make the model and its subsequent analysis substantially more complex.
We believe that our results, together with [38, 35], have set the stage for opening a new direction of research aimed at theoretical understanding of neurological disorders and associated therapeutic procedures by implementing the next-generation neural mass models. This is expected to provide a valuable complement to experimental studies in at least two ways. Firstly, it will allow for a qualitative insight into multistability which may be lacking or not be easily accessible to experiments; and secondly, it may help in developing effective strategies to control the switching dynamics, especially in terms of suppressing the pathological states. Concerning the research on the ketogenic diet, the next step could involve incorporating some other ingredients that may be related to its therapeutic effects. In particular, one may attempt to derive a “compartmental” model of local dynamics that could distinguish between the two different ATP production processes, namely the classical, glycolytic one near the membrane and the diet-related one unfolding in mitochondria. Also, one may attempt to circumvent the limitations of the current model associated with the assumed uniformity of ATP concentration over the population, aiming to accommodate for the neuron-specific ATP concentration variables that depend on the spike trains of individual neurons [69]. The microscopic system introduced in the latter way would likely show significant deviations from the current mean-field model, similar to those observed in [37, 40]. Finally, the immediate impact of ketone bodies on inhibition of excitatory synaptic channels could be taken into account as a factor modulating the synaptic efficacy. More generally, an interesting challenge would be to understand how the classical physical picture developed for activity of neuronal populations within the NGNM framework changes by incorporating the different forms of homeostatic plasticity [103], which involve negative feedback loops to constrain the network activity within the desired physiological limit. Also, as a next step, it would be important to provide at least a qualitative validation of our theoretical predictions through elaborately designed experiments, which could then lead to further development and refinement of the current model.